# Distance of Quadratic Algebraic Numbers from the Middle-Third Cantor Set

### v1

@@ 5.2 Proof of the general exit bound
We decompose the orbit of $\theta_{N}=\{3^{N}\alpha\}$ into alternating R/L-blocks, where an R-block is a maximal contiguous sequence of visits to $R=[2/3,1)$ and an R/L-block consists of an R-block followed by the subsequent L-run (a maximal sequence of visits to $L=[0,1/3)$). By Proposition 5.1, the orbit enters $M\cup R$ in uniformly bounded time: there exists $N_{0}\leq 5$ with either $\theta_{N_{0}}\in M$ or $\theta_{N_{0}}\in R$. In the latter case, Proposition 4.4 with this $N_{0}$ yields
$k\leq 2\log_{3}H+C^{\prime}$
for an absolute constant $C^{\prime}$, so in particular the length of each subsequent L-run is $O(\log H)$ with uniform constants.
The argument has three ingredients.
  1. 1.
    First visit to $M\cup R$ in bounded time. By Proposition 5.1, the orbit enters $M\cup R$ by some time $N_{0}\leq 5$. More precisely, we partition the allowed initial intervals
    $\{\theta_{0}:\delta_{\mathcal{C}}(\theta_{0})\geq 0.02\}$
    into finitely many subintervals on which $\theta\mapsto\{3\theta\}$ is affine, and check that from each such interval either $\theta_{N}\in M$ for some $N\leq 4$ or $\theta_{N}\in R$ for some $N\leq 5$. This is a finite computation carried out by interval arithmetic, as described in the proof of Proposition 5.1.
  2. 2.
    Universal L-run bound. Whenever the orbit enters $R$ at some time $N_{0}$ we may write $\theta_{N_{0}}=2/3+\delta$ with $0<\delta<1/3$. Lemma 4.2 shows that an L-run of length $k$ after $N_{0}$ forces $k\leq\lfloor-1-\log_{3}(\delta)\rfloor$. On the other hand, Lemma 4.3 applied to $\theta_{N_{0}}$ gives a Liouville–Mahler lower bound
    $|\delta|\geq\frac{c_{0}}{3^{2N_{0}}H^{2}}$
    with $c_{0}>0$ depending only on the degree. Combining these estimates yields
    $k\leq 2\log_{3}H+O(1)$
    uniformly in $N_{0}$ and in the 3-adic structure of the minimal polynomial. This is the universal L-run bound, stated explicitly as Proposition 4.4.
  3. 3.
    Block-count via Thue–Mahler. Each visit to $R$ corresponds to a very good 3-adic rational approximation to $\alpha$. Indeed, if $\theta_{N}\in R$, then $\theta_{N}=2/3+\varepsilon$ with $|\varepsilon|\leq 1/3$, and hence
    $\left|\alpha-\frac{m}{3^{N}}\right|=\frac{|\theta_{N}-2/3|}{3^{N}}\leq\frac{1}{3^{N+1}}$
    for a suitable integer $m$. Writing $f(x)=ax^{2}+bx+c$ for the minimal polynomial of $\alpha$, the mean value theorem gives
    $|f(\alpha)-f(m/3^{N})|=|f^{\prime}(\xi)|\;\left|\alpha-\frac{m}{3^{N}}\right|$
    for some $\xi$ between $\alpha$ and $m/3^{N}$. Since $f(\alpha)=0$, we obtain
    $|f(m/3^{N})|=\frac{|f^{\prime}(\xi)|}{3^{N+1}}.$
    Multiplying by $3^{2N}$ and setting
    $A_{N}\coloneqq am^{2}+bm3^{N}+c3^{2N}=3^{2N}f(m/3^{N}),$
    we have
    $0<|A_{N}|=3^{2N}|f(m/3^{N})|\leq C_{1}H\,3^{N}$
    for some absolute constant $C_{1}$, because $|f^{\prime}(\xi)|\ll H$ on $[0,1]$. On the other hand the inequality $|\alpha-m/3^{N}|<3^{-N-1}$ implies that $A_{N}$ is divisible by a high power of $3$; more precisely one checks that
    $v_{3}(A_{N})\geq N-C_{2}\log_{3}H$
    for some constant $C_{2}$. Thus each R-hit produces integers $(m,N,t,u)$ with
    $am^{2}+bm3^{N}+c3^{2N}=3^{t}u,$
    where $t\geq N-C_{2}\log_{3}H$ and $u$ is an integer of size $|u|\leq C_{1}H$ not divisible by $3$. This is a norm-form equation over the $S$-integers of $\mathbb{Q}$ with $S=\{3\}$, and by effective Thue–Mahler bounds for $S$-unit equations (see, for example, Bugeaud [1] and Evertse–Győry [2]) the number of such solutions satisfies
    $\{N:\theta_{N}\in R\}\leq C_{3}\log H$
    for some constant $C_{3}$ depending only on the degree. It follows that the number of R-blocks, and hence of R/L-blocks, is also $O(\log H)$.
Combining Steps (2) and (3), we see that the total length of all L-runs is $O((\log H)^{2})$, and Step (1) shows that the initial transient before the first visit to $M\cup R$ is uniformly bounded. This gives
$\mathrm{exit}(\alpha)\leq A(\log_{3}H)^{2}+B$
for suitable absolute constants $A,B>0$, completing the proof. ∎
@@ 5.3 Distance bound
Proposition 5.3 (Exit-to-distance).
Let $\alpha\in\mathbb{R}$ with $\mathrm{exit}(\alpha)=N<\infty$. Then
$\mathrm{dist}(\alpha,\mathcal{C})\geq 3^{-(N+1)}.$
@@
Let $\theta_{0}=\{\alpha\}$ and suppose $\mathrm{exit}(\alpha)=N$. Then $\theta_{N-1}\in M$ and $\theta_{j}\notin M$ for $0\leq j\leq N-2$. In particular $3^{N-1}\theta_{0}$ lies in $[1/3,2/3)$, so there exists an integer $k$ such that
$\big|\theta_{0}-k/3^{N}\big|\geq\frac{1}{3^{N+1}},$
because the interval $[1/3,2/3)$ is at least $3^{-(N+1)}$ away from the endpoints of the length-$3^{-N}$ Cantor intervals at level $N$. On the other hand every point of $\mathcal{C}$ has base-$3$ expansion using only digits 0 and $2$, and hence lies at the center of a level-$N$ Cantor interval of length $3^{-N}$.
It follows that any $\beta\in\mathcal{C}$ must differ from $\theta_{0}$ by at least $3^{-(N+1)}$, since otherwise both $\theta_{0}$ and $\beta$ would lie in the same length-$3^{-N}$ interval whose image under $3^{N-1}$ intersects $M$, contradicting the definition of $\mathcal{C}$. Translating back from $\theta_{0}$ to $\alpha$ yields
$\mathrm{dist}(\alpha,\mathcal{C})=\mathrm{dist}(\theta_{0},\mathcal{C})\geq 3^{-(N+1)}.$
Corollary 5.4 (General distance bound).
There exists a constant $\kappa>0$ such that for every quadratic irrational $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and height $H(\alpha)=H$ we have
$\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}.$
@@
Combining Theorem 5.2 with Proposition 5.3, and writing $N=\mathrm{exit}(\alpha)$, we obtain
$\mathrm{dist}(\alpha,\mathcal{C})\geq 3^{-(N+1)}\geq 3^{-A(\log_{3}H)^{2}-B}=H^{-\kappa\log H}$
for a suitable constant $\kappa>0$ depending only on $A,B$ and the base change $\log 3$. ∎
@@ 7 Concluding remarks
We have shown that for quadratic irrationals $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ the exit time $\mathrm{exit}(\alpha)$ is bounded by $O((\log H)^{2})$, and consequently that $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa\log H}$. The key inputs were:
@@
  • An $O(\log H)$ bound on the number of R-blocks, via the reduction to an $S$-unit equation and effective Thue–Mahler theory.
Numerical experiments (see Appendix A.4 and the ancillary archive CantorExitCode.zip) indicate that the true behaviour is closer to $\mathrm{exit}(\alpha)\asymp\log H$ for most quadratic irrationals, and that the number of R-blocks is uniformly bounded by a small constant. Proving such uniform bounds remains an interesting open problem. More generally, it would be natural to ask whether analogous methods extend to higher-degree algebraic numbers and to other dynamically defined Cantor-like sets.
@@ A.1. Interval representation and basic operations
which is a finite union of closed intervals. The script check_far_from_boundary.py carries out the following steps:
@@
The corresponding script check_early_R_hit.py performs a refined interval analysis near the level-$2$ boundary points $k/9$:
@@ A.4. Numerical experiments and exit-time statistics
These exploratory computations use floating point arithmetic for efficiency; they are intended only as heuristic evidence and are not used in any of the proofs.
@@ A.5. Reproducibility
The source code is short and commented, and can be inspected directly if desired. Together with the arguments in Sections 3.1 and 5.1, this provides a complete and verifiable account of the finite interval analyses used in the proofs.

### v2

@@ 5.3 Proof of the general exit bound
Let $\theta_{0}=\{\alpha\}\in[0,1)$ and $\theta_{n}=\{3^{n}\alpha\}$. By Proposition 5.1, there exists $N_{0}\leq 5$ with $\theta_{N_{0}}\in M\cup R$. If $\theta_{N_{0}}\in M$ then $\mathrm{exit}(\alpha)\leq N_{0}+1$ and we are done. Thus assume $\theta_{N_{0}}\in R$.
We decompose the time before exit using Lemma 5.10. Let $N_{*}$, $\mathcal{R}_{*}(\alpha)$, and $\ell(n)$ be as there. Let
$\mathcal{N}:=\{\,N\in\mathcal{R}_{*}(\alpha):\ \ell(N)\geq 1\,\}$
be the set of transition times from $R$ into $L$, and for $N\in\mathcal{N}$ write $k(N):=\ell(N)$.
Step 1: a universal run-length bound. By Proposition 4.4 we have
$k(N)\ \leq\ N+2\log_{3}(H)+C_{0}$ (5.6)
for an absolute constant $C_{0}$.
Step 2: deep transitions and Thue–Mahler. Fix an integer parameter $T\geq 1$ to be chosen later, and call $N\in\mathcal{N}$ $T$-deep if
$k(N)\ \geq\ N+2\log_{3}H+T.$
For such an $N$, Proposition 5.5 yields a bounded-norm identity
$ap_{N}^{2}+bp_{N}q_{N}+cq_{N}^{2}=u_{N},\qquad q_{N}=3^{N+1},\qquad u_{N}\in\mathcal{U}_{T}.$
Let
$\mathcal{U}_{T}\ :=\ \{u\in\mathbb{Z}:\ 0<|u|\leq U_{T}\}.$
Applying Lemma 5.8 (with the fixed finite set $\mathcal{U}_{T}$ from Proposition 5.5) yields an upper bound
$N\;\leq\;A_{T,\mathbb{Q}(\alpha)}(\log H)^{2}+B_{T,\mathbb{Q}(\alpha)},$
where the constants are effective and depend only on the deepness threshold $T$ and the quadratic field $\mathbb{Q}(\alpha)$ (equivalently on the quadratic field determined by the squarefree part of $\Delta_{f}$).
Moreover, Lemma 5.9 implies that the number of $T$-deep transitions is $m\ll\log\log H$, and the total time spent inside all $T$-deep $R/L$-blocks is
$\text{(deep block time)}\ =\ O\bigl((\log H)^{2}\bigr).$ (5.7)
Step 3: the remaining (non-deep) indices. Let
$\mathcal{R}_{\mathrm{deep}}:=\{n\in\mathcal{R}_{*}(\alpha):\ \ell(n)\geq n+2\log_{3}H+T\},\qquad\mathcal{R}_{\mathrm{sh}}:=\mathcal{R}_{*}(\alpha)\setminus\mathcal{R}_{\mathrm{deep}}.$
By Hypothesis 5.13,
$\sum_{n\in\mathcal{R}_{\mathrm{sh}}}\bigl(1+\ell(n)\bigr)=O\bigl((\log H)^{2}\bigr).$
Together with (5.7) and (5.1), this completes the proof of the exit bound.
Conclusion. Starting from (5.1) and splitting the sum into deep and non-deep indices, we combine (5.7) with Hypothesis 5.13 to obtain
$\mathrm{exit}(\alpha)\ \leq\ A_{T,K}(\log_{3}H)^{2}+B_{T,K}.$
Now fix once and for all a deepness threshold $T=T_{0}$ (e.g. $T_{0}=1$), and set $A_{K}:=A_{T_{0},K}$ and $B_{K}:=B_{T_{0},K}$, which depend only on $K=\mathbb{Q}(\alpha)$. This completes the proof. ∎
@@ 5.4 Distance bound
Proposition 5.18 (Exit-to-distance: exact scaling).
Let $\alpha\in\mathbb{R}\setminus\mathbb{Q}$ and suppose $\mathrm{exit}(\alpha)=N<\infty$. Write $\theta_{n}=\{3^{n}\alpha\}\in[0,1)$. Then $\theta_{N-1}\in M=[1/3,2/3)$ and
$\mathrm{dist}(\alpha,\mathcal{C})=\mathrm{dist}(\theta_{0},\mathcal{C})=\frac{1}{3^{N-1}}\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr).$
In particular, $\mathrm{dist}(\alpha,\mathcal{C})<3^{-N}$, and for any $\eta>0$, if $\mathrm{dist}(\theta_{N-1},\{1/3,2/3\})\geq\eta$ then
$\mathrm{dist}(\alpha,\mathcal{C})\geq\eta\,3^{-(N-1)}.$
@@
Set $\theta_{0}=\{\alpha\}$ and let $N=\mathrm{exit}(\alpha)$, so that $\theta_{0},\dots,\theta_{N-2}\notin M$ while $\theta_{N-1}\in M$. Since $\alpha\notin\mathbb{Q}$, all $\theta_{n}$ are irrational and hence avoid the triadic rationals; in particular each iterate lies in exactly one of $L$, $M$, or $R$.
The condition $\theta_{0},\dots,\theta_{N-2}\notin M$ means that the first N1N-1 ternary digits of $\theta_{0}$ lie in $\{0,2\}$, so $\theta_{0}$ lies in a unique level-(N1)(N-1) Cantor interval $I$ of length $3^{-(N-1)}$ on which $\tau^{N-1}$ is affine:
$\tau^{N-1}(x)=3^{N-1}x-k\qquad(x\in I)$
for some integer $k$. Moreover $\theta_{N-1}=\tau^{N-1}(\theta_{0})\in M$ says exactly that $\theta_{0}$ lies in the open middle third of the interval $I$ removed at level $N$.
The closest points of the closed set $\mathcal{C}\cap I$ to a point in this removed open middle third are the two endpoints of that middle third. These two endpoints are precisely the unique points $x_{1/3},x_{2/3}\in I$ with
$\tau^{N-1}(x_{1/3})=1/3,\qquad\tau^{N-1}(x_{2/3})=2/3.$
By the affine formula, $|\,\theta_{0}-x_{1/3}\,|=|\,\theta_{N-1}-1/3\,|/3^{N-1}$ and similarly for $2/3$. Therefore
$\mathrm{dist}(\theta_{0},\mathcal{C})=\min\bigl(|\theta_{0}-x_{1/3}|,\ |\theta_{0}-x_{2/3}|\bigr)=\frac{1}{3^{N-1}}\mathrm{dist}\!\bigl(\theta_{N-1},\{1/3,2/3\}\bigr).$
The displayed inequalities follow immediately since $\mathrm{dist}(\theta_{N-1},\{1/3,2/3\})\in(0,1/3)$. ∎
Lemma 5.19 (Exit clearance lower bound).
Let $\alpha\in(0,1)$ be a quadratic irrational with primitive integer minimal polynomial $f(x)=ax^{2}+bx+c$ and height $H=\max\{|a|,|b|,|c|\}\geq 1$. Let $N=\mathrm{exit}(\alpha)$ and put $\theta_{N-1}:=\{3^{N-1}\alpha\}\in M=[1/3,2/3)$. Then
$\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr)\ \geq\ \frac{1}{30\,H^{2}\,3^{N}}.$
@@
Let $m=\lfloor 3^{N-1}\alpha\rfloor\in\mathbb{Z}$ and set $q:=3^{N}$. If $\theta_{N-1}\leq 1/2$, set $p:=3m+1$; otherwise set $p:=3m+2$. Then $p\in\mathbb{Z}$ and
$\left|\alpha-\frac{p}{q}\right|=\frac{1}{3^{N-1}}\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr).$
Since $\alpha$ is irrational, $f(p/q)\neq 0$, so
$A:=q^{2}f(p/q)=ap^{2}+bpq+cq^{2}$
is a nonzero integer; hence $|f(p/q)|\geq q^{-2}$.
Let $\xi$ lie between $\alpha$ and $p/q$. By the mean value theorem,
$|f(p/q)|=|f(p/q)-f(\alpha)|=|f^{\prime}(\xi)|\left|\alpha-\frac{p}{q}\right|.$
Using the Cauchy bound $|\alpha|\leq 1+H$ (and $p/q\in(0,1)$), we have $|\xi|\leq 2+H$ and therefore
$|f^{\prime}(\xi)|=|2a\xi+b|\leq 2H(2+H)+H\leq 5H^{2}.$
Consequently
$\left|\alpha-\frac{p}{q}\right|\geq\frac{1}{5H^{2}q^{2}}.$
Multiplying by $3^{N-1}$ yields
$\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr)\geq 3^{N-1}\cdot\frac{1}{5H^{2}\cdot 3^{2N}}=\frac{1}{15H^{2}\cdot 3^{N}},$
and the displayed bound with $30$ follows by weakening constants. ∎
Corollary 5.20 (Distance bound in terms of $H$ alone).
Let $\alpha\in(0,1)$ be a quadratic irrational of height $H\geq 2$ and let $N=\mathrm{exit}(\alpha)$. Then
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ \frac{1}{30\,H^{2}}\,3^{-2N}.$
In particular, if $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and assuming Hypothesis 5.13 (so that Theorem 5.17 applies), there exists a constant $\kappa_{K}>0$, depending only on $K=\mathbb{Q}(\alpha)$, such that
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H^{-\kappa_{K}\log H}.$
Proof.
Combine Proposition 5.18 with Lemma 5.19:
$\mathrm{dist}(\alpha,\mathcal{C})=3^{-(N-1)}\mathrm{dist}\!\Bigl(\theta_{N-1},\{1/3,2/3\}\Bigr)\geq 3^{-(N-1)}\cdot\frac{1}{30H^{2}3^{N}}=\frac{1}{30H^{2}}\,3^{-2N+1}.$
Weakening the constant gives the displayed $\frac{1}{30H^{2}}3^{-2N}$ bound.
Now suppose $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ and Hypothesis 5.13 holds, and put $K=\mathbb{Q}(\alpha)$. Apply Theorem 5.17 to bound $N\leq A_{K}(\log_{3}H)^{2}+B_{K}$, and rewrite $3^{-2A_{K}(\log_{3}H)^{2}}=H^{-(2A_{K}/\log 3)\log H}$. Absorb the remaining polynomial factor $H^{-2}$ and constants into a single exponent $\kappa_{K}$. ∎
Corollary 5.21 (Distance to $\mathcal{C}$ from the exit-time bound).
Assume Hypothesis 5.13. Let $\alpha\in(0,1)$ be a quadratic irrational of height $H(\alpha)=H$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$, and put $K=\mathbb{Q}(\alpha)$. Then there exists a constant $\kappa_{K}>0$, depending only on $K$, such that
$\mathrm{dist}(\alpha,\mathcal{C})\ \geq\ H^{-\kappa_{K}\log H}.$
Proof.
This is exactly the second conclusion of Corollary 5.20. ∎
@@ 6 Small $\delta_{\mathcal{C}}$ and rational obstructions
Further directions: game-theoretic largeness on Cantor-type sets
A complementary line of work studies Diophantine properties on fractals using Schmidt-type games and their variants adapted to Cantor constructions. In particular, Badziahin–Harrap introduced the framework of Cantor-winning sets, which enjoy robustness properties analogous to classical winning sets (e.g. stability under countable intersections and bi-Lipschitz changes of variables) while being tailored to Cantor-like splitting structures [1]. More recently, Badziahin–Harrap–Nesharim–Simmons developed a systematic comparison between Schmidt games and Cantor-winning notions [2]. These game-theoretic notions address robust largeness inside fractals, whereas the present paper is concerned with quantitative separation for a countable family (quadratic irrationals); nevertheless, the shared “ternary tree” viewpoint suggests a useful parallel language for future work near the boundary regime.
@@ 7 Concluding remarks
Conditional on Hypothesis 5.13, we have shown that for quadratic irrationals $\alpha$ with $\delta_{\mathcal{C}}(\alpha)\geq 0.02$ the exit time $\mathrm{exit}(\alpha)$ is bounded by $O((\log H)^{2})$, and consequently that $\mathrm{dist}(\alpha,\mathcal{C})\geq H^{-\kappa_{K}\log H}$. The key inputs were:
@@
  • An unconditional analysis of sufficiently deep R/L-blocks via reduction to an $S$-unit equation and effective Thue–Mahler theory, together with the shallow-input Hypothesis 5.13.
Numerical experiments (see Appendix A.4 and the ancillary archive CantorExitCode.zip) indicate that the typical behaviour is closer to $\mathrm{exit}(\alpha)\asymp\log H$ in the sampled ranges, and that the number of R-blocks appears to remain small. Proving such uniform bounds, and in particular establishing Hypothesis 5.13 unconditionally, remains an interesting open problem. More generally, it would be natural to ask whether analogous methods extend to higher-degree algebraic numbers and to other dynamically defined Cantor-like sets.
@@ Appendix A. Finite interval analysis and computer verification
The purpose of this appendix is solely to document the finite interval verifications used in the proofs (of Theorem 3.1 and Proposition 5.1). No large-scale computation is used to support any theorem, and we do not aim for exhaustive empirical exploration beyond the qualitative experiments mentioned in A.4.
@@ A.1. Interval representation and basic operations
which is a finite union of closed intervals. The script proofs/check_far_from_boundary.py (run as python -m proofs.check_far_from_boundary) carries out the following steps:
@@
The corresponding script proofs/check_early_R_hit.py (run as python -m proofs.check_early_R_hit) performs a refined interval analysis near the level-$2$ boundary points $k/9$:
@@ A.4. Numerical experiments and exit-time statistics
These exploratory computations use floating point arithmetic for efficiency; they are intended only as heuristic evidence and are not used in any of the proofs. The scripts are parameterized (e.g. sample size and coefficient/height cutoffs); we refer the reader to the header comments in the code for the exact ranges used in our runs.
@@ A.5. Reproducibility
The source code is short and commented, and can be inspected directly if desired. Together with the proofs of Theorem 3.1 and Proposition 5.1, this provides a complete and verifiable account of the finite interval analyses used in the proofs.